A wire carrying a 25-A current passes between the poles of a strong magnet that is perpendicular to its field and experiences a 5.96-N force on the 3.86 cm of wire in the field. What is the average field strength?


1
Expert's answer
2021-05-25T16:49:58-0400

The magnitude of the force on a wire of length l carrying a current I and making an angle "\theta" with a magnetic field of density B given by :

"|\overrightarrow{AB}| = B.I.l.sin(\theta)"


Rearranging to get "\overrightarrow{B}"


"B = \dfrac{|\overrightarrow{F}| }{Ilsin\theta}\\
\space\space\space\space = \dfrac{5.96}{25\times(3.86\times10^{-2})\times sin \dfrac{\pi}{2}}\\
\space\space\space\space= \dfrac{5.96}{0.965}\\"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space = 6.167T"

answer B= 6.167T